You are on an interstellar mission from the Earth to the 8.7 light-years distant star Sirius. Your
spaceship can travel with 70% the speed of light and has a cylindrical shape with a diameter of
6 m at the front surface and a length of 25 m. You have to cross the interstellar medium with an
approximated density of 1 hydrogen atom/m3
.
(a) Calculate the time it takes your spaceship to reach Sirius.
(b) Determine the mass of interstellar gas that collides with your spaceship during the mission.
1
Expert's answer
2020-06-18T11:36:25-0400

a)


"t=\frac{8.7}{0.7}=12.4\ years"

b)


"M=m_pN=m_p\rho V=m_p\rho (0.25\pi d^2)s"

"M=(1.67\cdot10^{-27})(1)(0.25\pi (6)^2)(8.7\cdot9.46\cdot10^{15})\\=3.89\cdot10^{-9}\ kg"